40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Mar 6 12:34:18 2021, gautam, Update, LSC, Sensing matrix settings messed with sensMat.png
    Reply  Sun Mar 7 19:56:27 2021, Anchal, Update, LSC, Sensing matrix settings messed with 
Message ID: 15876     Entry time: Sun Mar 7 19:56:27 2021     In reply to: 15874
Author: Anchal 
Type: Update 
Category: LSC 
Subject: Sensing matrix settings messed with 

I understand this mst be frustrating for you. But we did not change these settings, knowingly atleast. We have documented all the things we did there. The only thing I can think of which could possibly change any of those channels are the scripts that we ran that are mentioned and the burt restore that we did on all channels (which wasn't really necessary). We promise to be more vigilant of changes that occur when we are present in future.

Quote:

To my dismay, I found today that somebody had changed the oscillator frequencies for the sensing matrix infrastructure we have. The change happened 2 days and 2 hours ago (I write this at ~1230 on Saturday, 3/6), i.e. ~1030am on Thursday. According to the elog, this is when Anchal and Paco were working on the interferometer, but I can find no mention of these settings being changed. Not cool guys 😒 .

This was relatively easy to track down but I don't know what else may have been messed with. I don't understand how anything that was documented in the elog can lead to this weird doubling of the frequencies.

I have now restored the correct settings. The "sensing matrix" I posted last night is obviously useless.

 

ELOG V3.1.3-